a ball which is thrown vertically upwards and reaches the roof of a house 100 m high. At the moment this ball is thrown vertically upward , antoher ball is dropped from rest evrtically downward from the roof of the house. at which height will the balls paas each other . ( g =9.8 m/s^2 )

Grade:12

2 Answers

Ruchi Thakur
9 years ago
For the first ball
taking up as positive and don negative
v=u+at
for total tme of flight, T
-u=u-gT or T = 2u/g
displacement, s=ut+o.5t2 
at the highst point, 100= ut – 0.5gt2 , becoz t=0.5T , thus 100= u2/g – o.5g (u/g)2
 or u2 = 200g ,,,
,u =10(2g)½
for the trajectory of first ball y= 10(2g)1/2t -5t2 
bus ab mai thak gaye
Ruchi Thakur
9 years ago
for the trajectory of 2nd ball
displacement, s= 0 -0.5gt2
position, y = 100 -0.5 gt2
on comparison of the 2 eqns, y = 100 – 0.5gt= 10(2g)1/2t-0.5gt2
or  t= 10(2g)-1/2  thus t2= 50/g...
thus y = 100 – 0.5gt2 = 100- 25= 75 m from the initial point of the first ball.